OpenStudy (anonymous):

A particle of mass m = 3 kg is released from point A on a frictionless track at height hA = 6.28 m. The acceleration of gravity is 9.8 m/s 2 . Determine the particle’s speed at point B at height hB = 2.2 m. Answer in units of m/s

OpenStudy (anonymous):

in this case, GPE = gravitational potential energy the amount of kinetic energy = GPE_6.28m - GPE_2.2m = 3(6.28)(9.8) - 3(2.2)(9.8) = 120.0 KE = 0.5mv^2 = 120.0 therefore, speed = 120.0/(0.5)(3) = 40m/s